What two numbers have the sum of 9 and the product of 20?
Answer by bucky(2189)

Call the two numbers x and y
.
Since their sum is 9 we write this in equation form as:
.
x + y = 9
.
And since their product is 20 we write this in equation form as:
.
x*y = 20
.
We now have two equations. One way to solve an equation set such at this is by using substitution.
.
Solve the first equation regarding the sum for one of the variables in terms of the other.
You can do this by subtracting y from both sides to get:
.
x + y - y = 9 - y
.
Perform the subtraction on the left side and the equation reduces to:
.
x = 9 - y
.
Then in the second equation we can substitute 9 - y for x since x is equal to 9 - y.
After making this substitution the second equation becomes:
.
(9 - y)*y = 20
.
Doing the distributed multiplication on the left side (multiplying the y by each of the terms
in the set of parentheses) results in the equation becoming:
.
9y - y^2 = 20
.
rearrange the terms on the left side in descending powers of y to get:
.
-y^2 + 9y = 20
.
Subtract 20 from both sides and the equation becomes:
.
-y^2 + 9y - 20 = 0
.
Finally, to get things in a little more standard form, multiply all the terms on both sides
by -1 to make the y^2 term be positive. After this multiplication, the equation becomes:
.
y^2 - 9y + 20 = 0
.
This factors into:
.
(y - 5)*(y - 4) = 0
.
This equation will be true if either one of the factors equals zero, because a multiplication
by zero on the left side will result in that side becoming zero and therefore it equals the
right side.
.
So set y - 5 = to 0. Solve for y by adding 5 to both sides to get y = 5.
.
Then set y - 4 equal to 0. Solve for y by adding 4 to both sides to get y = 4.
.
So we have two possibilities ... either y = 5 or y = 4.
.
Now go to the equation that said x + y = 9. If y is equal to 5 this equation becomes
x + 5 = 9 and when you subtract 5 from both sides this tells you that x = 4. However,
if y is equal to 4 this equation becomes x + 4 = 9 and when you subtract 4 from both sides
you get that x = 5.
.
So the answer is that the two numbers are 4 and 5. Their sum is 9 and their product is 20.
.
Hope this helps you to see your way through this problem.